OpenStudy (thatonegirl_):

For a company manufacturing iPADs the cost and revenue equations (in dollars) are given by... (below) where x is the weekly production output of iPADs. If production is increasing at a rate of 4 iPADs per week when 360 iPADs are produced, find the rate of increase in profit.

OpenStudy (thatonegirl_):

\[C(x)=200+30x\] \[R(x)=300x-\frac{ x^2 }{ 40 }\]

OpenStudy (solomonzelman):

What is C(x) and R(x) ?

OpenStudy (thatonegirl_):

I am assuming cost and revenue

OpenStudy (solomonzelman):

The profit is revenue-cost. (Right?) (I will denote profit by \(P(x)\).) \(P(x)=R(x)-C(x)\)

OpenStudy (thatonegirl_):

Yup looks good

OpenStudy (solomonzelman):

So, subtract, and that would be the function that models your profit.

OpenStudy (solomonzelman):

Then, if you want to get the rate of increase in profit, just differentiate (once).

OpenStudy (thatonegirl_):

Okay gimme a min

OpenStudy (thatonegirl_):

Should i be using the quotient rule?

OpenStudy (solomonzelman):

First, tell me what is your \(P(x)\) ?

OpenStudy (solomonzelman):

(no, I don't think you should)

OpenStudy (thatonegirl_):

\[\frac{ -x^2+1080x-8000 }{ 40 }\]

OpenStudy (thatonegirl_):

i combined them all. Could i just multiply by 40 to get rid of the denominator?

OpenStudy (solomonzelman):

\(\displaystyle P(x)=270x-\frac{x^2}{40}-200\)

OpenStudy (solomonzelman):

You don't need to combine.

OpenStudy (solomonzelman):

Just differentiate with the Power Rule, term by term.

OpenStudy (thatonegirl_):

is this with respect to time or to x?

OpenStudy (solomonzelman):

Oh, with respect to x, of course.

OpenStudy (thatonegirl_):

oh okay

OpenStudy (thatonegirl_):

hmm so how do u take the derivative of x^2/40?

OpenStudy (solomonzelman):

How do you differentiate \(\displaystyle ax^2\), or \(\displaystyle \frac{1}{40}x^2\) ?

OpenStudy (thatonegirl_):

ohh ty lol

OpenStudy (thatonegirl_):

okay so \[P'(x)=270-\frac{ 1 }{ 20 }x\]

OpenStudy (solomonzelman):

Yes.

OpenStudy (thatonegirl_):

Okay so then what?

OpenStudy (solomonzelman):

I think we found the rate of increase in profit, haven't we?

OpenStudy (thatonegirl_):

my answers choices are $1007per week $1006per week $1008per week $1010per week $1009 per week

OpenStudy (solomonzelman):

I might have gone wrong somewhere:(

OpenStudy (thatonegirl_):

If it helps at all we are learning about derivatives with respect to time >.<

OpenStudy (thatonegirl_):

Idk the fancy name for it lol oops

OpenStudy (solomonzelman):

Apparently the rate of change of the profit is there, but I don't know how to relate that to the verbal part of the problem.

OpenStudy (solomonzelman):

(I am not a very patient reader, perhaps. Sorry.)

OpenStudy (thatonegirl_):

Ah okay thanks anyways! I appreciate your time and effort!! :)